# Astronomical Cardiology: A Search For Heartbeat Stars Using Gaia and TESS

### v1

@@ Abstract
Heartbeat stars are a subclass of binary stars with short periods, high eccentricities, and phase-folded light curves that resemble an electrocardiogram. We start from the Gaia catalogs of spectroscopic binaries and use TESS photometry to identify 112 new heartbeat star systems. We fit their phase-folded light curves with an analytic model to measure their orbital periods, eccentricities, inclinations, and arguments of periastron. We then compare these orbital parameters to the Gaia spectroscopic orbital solution. Our periods and eccentricities are consistent with the Gaia solutions for 85 of the single-line spectroscopic binaries but only 20 of the double-line spectroscopic binaries. For the two double-line spectroscopic binary heartbeat stars with consistent orbits, we combine the TESS phase-folded light curve and the Gaia velocity semi-amplitudes to measure the stellar masses and radii with PHOEBE. In a statistical analysis of the HB population, we find that non-giant heartbeat stars have evolved off the main sequence and that their fractional abundance rises rapidly with effective temperature.
@@
Heartbeat stars (HB) are a relatively rare subclass of detached binary stars with short periods ( days) and high eccentricities (). In these binaries, the stars are tidally deformed near pericenter, causing periodic brightness variations. HBs are easily identifiable because their phase-folded light curves resemble an electrocardiogram (Thompson et al., 2012), and a simple analytic model (Kumar et al., 1995) can be used to estimate the properties of the orbit. The tides can also produce tidally excited oscillations (TEOs), waning pulsations of the stellar material that is observed between pericenter passages.
@@ 2. Heartbeat Search
We start from the 181,529 SB1s and 5,376 SB2s (Gaia Collaboration et al., 2023) in Gaia DR3. We used phase-folded light curves from the TESS Quick-Look Pipeline (QLP, Huang et al., 2020) for all of the Gaia SB1s and SB2s. The typical Gaia spectroscopic binary has four TESS sectors of observations.
@@ 2.1 SB2 Modeling
Since there are only 5,376 Gaia SB2 targets, we identified HB candidates through visual inspection of the TESS phase-folded light curves. For all targets, we run a Lomb-Scargle periodogram (LS, Lomb, 1976; Scargle, 1982) on each TESS sector independently. We visually inspected the phase-folded light curves and identified 10 HB candidates. We then use phase dispersion minimization (PDM, Stellingwerf, 1978) to refine the orbital period. Before modeling the phase-folded light curve, we select a single TESS sector for each target and remove long-term trends in the data, if necessary. For example, TIC 265056364 has a continuous decrease in the flux over the course of the TESS sector, which could be due to variable blended light or systematic trends. We remove these signals with a linear fit.
@@ 2.3 The Final Sample
Fig. 5 shows the orbital periods and eccentricities of the 112 HBs detected in the Gaia SB1 and SB2 samples. Of these 112 HBs, four were found by Solanki et al. (2025) (TIC 252588526, TIC 271554988, TIC 344586348, and TIC 370209445). The orbital periods range from 1.51 to 12.17 days, with a median of 4.36 days. The eccentricities range from 0.10 to 0.57 (TIC 42430804), with a median of 0.20. All of our targets fall below the period-eccentricity envelope expected from tidal circularization (Mazeh, 2008). Fig. 5 also shows the HBs from Wrona et al. (2022a, b), Li et al. (2023), and Solanki et al. (2025). As compared to the HBs identified in Wrona et al. (2022a, b), our systems have much shorter periods than their median value of 265.3 days. The median eccentricities are similar (0.24 for OGLE). Our period cut will exclude giants while the OGLE magnitude limits favour giants given the distance to the Magellanic Clouds.
@@ 3. Gaia Comparison
Fig. 9 shows the posteriors for TIC 98552498. The mass and radius of the primary are (2.61 0.27) and (1.41 0.40) , while the mass and radius of the secondary are (2.53 0.29) and (2.85 0.11) . Fig. 10 shows the posteriors for TIC 76094846. The mass and radius of the primary are (8.19 1.19) and (5.14 0.18) , while the mass and radius of the secondary are (8.11 1.39) and (4.97 0.47) . Both systems are consistent with masses and radii of massive stars measured from detached eclipsing binaries (Torres et al., 2010).
@@ 4. Discussion
To examine the second point, we computed the fraction of HBs as a function of color. There are two reasons the HB fraction might drop as we examine stars lower on the MS. First, the redder, lower mass stars have longer and longer evolutionary time scales, which will lead to lower amplitudes (MacLeod & Loeb, 2025). Second, stars below the Kraft break ( K or for the isochrone in Fig. 11, Kraft, 1967; Schatzman, 1962) have convective envelopes and so should have shorter damping time scales. Both of these effects will lead to lower HB amplitudes. We have some ability to distinguish the two issues because the highest mass stars below the Kraft break (Kraft, 1967; Schatzman, 1962) can be old enough to have started to evolve off the MS. We computed the fractions of SB1 and SB2 HBs as a function of color again, but along lines parallel to the one labeled "Kraft break" in Fig. 11 in mag bins. We report the median fractions and 90 confidence uncertainties or upper limits.
The results shown in Fig. 12 are striking. The fraction of HBs drops very rapidly for redder, cooler stars starting above the Kraft break (Kraft, 1967; Schatzman, 1962) and then may be flattening for still cooler stars. The SB2 fractions and limits are in agreement with those for the SB1s. The absolute normalizations of the SB1 fractions change when we use the full SB1 sample for the normalization of the fraction, rather than just those with periods days, but the trends are unchanged. On the upper MS, HBs are not rare - over 1 of Gaia SB1 systems bluer than mag are HBs.
@@ Acknowledgments
JC thanks Tharindu Jayasinghe and the Time Domain Research Group for their guidance, advice, and support. JC also thanks Collin Christy for his assistance through the years. DMR acknowledges support from the OSU Presidential Fellowship. CSK, KZS, and DMR are supported by NSF grants AST-2307385 and 2407206. ASAS-SN is funded by Gordon and Betty Moore Foundation grants GBMF5490 and GBMF10501 and the Alfred P. Sloan Foundation grant G-2021-14192.

### v2

@@ Abstract
Heartbeat stars are a subclass of binary stars with short periods, high eccentricities, and phase-folded light curves that resemble an electrocardiogram. We start from the Gaia catalogs of spectroscopic binaries and use TESS photometry to identify 112 new heartbeat star systems. We fit their phase-folded light curves with an analytic model to measure their orbital periods, eccentricities, inclinations, and arguments of periastron. We then compare these orbital parameters to the Gaia spectroscopic orbital solution. Our periods and eccentricities are consistent with the Gaia solutions for 85 of the single-line spectroscopic binaries but only 20 of the double-line spectroscopic binaries. For the two double-line spectroscopic binary heartbeat stars with consistent orbits, we combine the TESS phase-folded light curve and the Gaia velocity semi-amplitudes to measure the stellar masses and radii with PHOEBE. In a statistical analysis of the heartbeat star population, we find that non-giant heartbeat stars have evolved off the main sequence and that the fraction of the systems that are heartbeat stars rises rapidly with effective temperature.
@@
Heartbeat stars (HB) are a relatively rare subclass of detached binary stars with short periods ( days) and high eccentricities (). In these binaries, the stars are tidally deformed near pericenter, causing periodic brightness variations. HBs are easily identifiable because their phase-folded light curves resemble an electrocardiogram (Thompson et al., 2012), and a simple analytic model (Kumar et al., 1995) can be used to estimate the properties of the orbit. The tides can also produce tidally excited oscillations (TEOs) of the star that are observed between pericenter passages.
@@ 2. Heartbeat Search
We start from the 181,529 SB1s and 5,376 SB2s (Gaia Collaboration et al., 2023) in Gaia DR3. We used phase-folded light curves for sectors 1-79 from the TESS Quick-Look Pipeline (QLP, Huang et al., 2020) for all of the Gaia SB1s and SB2s. The typical Gaia spectroscopic binary has four TESS sectors of observations.
@@ 2.1 SB2 Modeling
Since there are only 5,376 Gaia SB2 targets, we identified HB candidates through visual inspection of the TESS phase-folded light curves. For all targets, we run a Lomb-Scargle periodogram (LS, Lomb, 1976; Scargle, 1982) on each TESS sector independently. We visually inspected the phase-folded light curves and identified 10 HB candidates. We then use phase dispersion minimization (PDM, Stellingwerf, 1978) to refine the orbital period. Before modeling the phase-folded light curve, we select a single TESS sector for each target and remove long-term trends in the data, if necessary. For example, TIC 265056364 has a continuous decrease in flux over the course of the TESS sector, which could be due to variable blended light or systematic trends. We remove these signals with a linear fit.
@@ 2.3 The Final Sample
Fig. 5 shows the orbital periods and eccentricities of the 112 HBs detected in the Gaia SB1 and SB2 samples. Of these 112 HBs, four were found by Solanki et al. (2025) (TIC 252588526, TIC 271554988, TIC 344586348, and TIC 370209445). We found no matches to the other comparison samples used below. The orbital periods range from 1.51 to 12.17 days, with a median of 4.36 days. The eccentricities range from 0.10 to 0.57 (TIC 42430804), with a median of 0.20. All of our targets fall below the period-eccentricity envelope expected from tidal circularization (Mazeh, 2008). Fig. 5 also shows the HBs from Wrona et al. (2022a, b), Li et al. (2023), and Solanki et al. (2025). As compared to the HBs identified in Wrona et al. (2022a, b), our systems have much shorter periods than their median value of 265.3 days. The median eccentricities are similar (0.24 for OGLE). Our period cut will exclude giants while the OGLE magnitude limits favour giants given the distance to the Magellanic Clouds.
@@ 3. Gaia Comparison
Fig. 9 shows the posteriors for TIC 98552498. The mass and radius of the primary are (2.61 0.27) and (1.41 0.40) , while the mass and radius of the secondary are (2.53 0.29) and (2.85 0.11) . Fig. 10 shows the posteriors for TIC 76094846. The mass and radius of the primary are (8.19 1.19) and (5.14 0.18) , while the mass and radius of the secondary are (8.11 1.39) and (4.97 0.47) . Both systems are consistent with the masses and radii of massive stars measured from detached eclipsing binaries (Torres et al., 2010).
@@ 4. Discussion
To examine the second point, we computed the fraction of HBs as a function of color. There are two reasons the HB fraction might drop as we examine stars lower on the MS. First, the redder, lower mass stars have longer and longer evolutionary time scales, which will lead to lower amplitudes (MacLeod & Loeb, 2025). Second, stars below the Kraft break ( K or for the isochrone in Fig. 11, Kraft, 1967; Schatzman, 1962; Beyer & White, 2024) have convective envelopes and so should have shorter damping time scales. Both of these effects will lead to lower HB amplitudes. We have some ability to distinguish the two issues because the highest mass stars below the Kraft break (Kraft, 1967; Schatzman, 1962; Beyer & White, 2024) can be old enough to have started to evolve off the MS. We computed the fractions of SB1 and SB2 HBs as a function of color again, but along lines parallel to the one labeled "Kraft break" in Fig. 11 in mag bins. We report the median fractions and 90 confidence uncertainties or upper limits.
The results shown in Fig. 12 are striking. The fraction of HBs drops very rapidly for redder, cooler stars starting above the Kraft break (Kraft, 1967; Schatzman, 1962; Beyer & White, 2024) and then may be flattening for still cooler stars. The SB2 fractions and limits are in agreement with those for the SB1s. The absolute normalizations of the SB1 fractions change when we use the full SB1 sample for the normalization of the fraction, rather than just those with periods days, but the trends are unchanged. On the upper MS, HBs are not rare - over 1 of Gaia SB1 systems bluer than mag are HBs.
@@ Acknowledgments
JC thanks Tharindu Jayasinghe and the Time Domain Research Group for their guidance, advice, and support. JC also thanks Collin Christy for his assistance through the years. DMR acknowledges support from the OSU Presidential Fellowship. CSK, KZS, and DMR are supported by NSF grants AST-2307385 and 2407206. ASAS-SN is funded by Gordon and Betty Moore Foundation grants GBMF5490 and GBMF10501 and the Alfred P. Sloan Foundation grant G-2021-14192. Thanks to Dr. J. Tayar for a communication on an earlier version of this manuscript.